# In-context learning of evolving data streams with tabular foundational models

### v1

@@ 3 In-context stream learning
Frequentistic perspective. Despite contexts being sampled from a different distribution than the pretraining distribution, the asymptotic prediction error of ICL is optimal when the signal about the latent concept in each context example is larger than the error due to the distribution mismatch. As a more informative and representative input space is provided, the learning error decreases [96]. In this regard, one can also take a frequentistic perspective [99]. From a predictor’s variance standpoint, being a pre-tuned, but untrained predictor with many hyperparameters and multi-head attention, a LTM has extremely high sensitivity to individual training samples, which translates in an increased ability to choose submodels and vanishing variance. From a bias standpoint, hyperparameters are pre-tuned to be optimal for a set of tasks defined by the prior. If this prior has large enough support and does not concentrate too much away from the true hypothesis, one can guarantee that the PPD converges to a close approximation of the true predictive distribution. Consequently, whether the LTM predictor can learn at inference time depends on its structural properties, with the optimal LTM approximation being characterized by a Kullback-Leibler criterion [100, 99]. To allow for accurate approximation of the conditional class probabilities, one needs sufficiently complex LTM models and prior. Practically, a LTM is trained on simulated data sets. The larger these data sets are, the more complex the PPD is approximated. The training set size can therefore be understood as a regularizer on the expected complexity of the network. Thus, while LTM implementations ensure vanishing predictor’s variance, the bias vanishes only if the provided data is somewhat localized around the test features.
3.1 Context optimization on evolving data streams
Based on this intuition, the objective of in-context stream learning is to minimize bias in model predictions when faced with new data that drifts over time. This problem occurs in situations where the LTM $f_{\theta}$, is applied on-the-fly to a continuous stream of real-world data. The data stream is assumed to be partitioned sequentially into non-overlapping segments, called concepts, such as $C_{i}$ and $C_{i+1}$, each corresponding to distinct joint data distributions, $p_{i}(x,y)$ and $p_{i+1}(x,y)$, respectively. Within each concept, the data distribution is assumed static, however, as the data stream progresses, drift between consecutive concepts can introduce discrepancies between past and future data. For instance, if concept $C_{i}$ has a skewed label distribution $p_{i}(y)$, and the following concept $C_{i+1}$ has a uniform label distribution $p_{i+1}(y)$, then the predictive model must adapt to account for this shift in label distribution [81]. The goal is to select an optimal localized context $l$, composed of $(x_{1},y_{1},\ldots,x_{n},y_{n},x_{\text{query}})$, which helps $f_{\theta}$ approximate the correct output $y_{\text{query}}$ even for cases in the tail regions of these drifting distributions. In this case, the model’s expected prediction error for a new sample $(x_{\text{query}},y_{\text{query}})$ is defined by:
$\begin{split}EPE_{f_{\theta}}(x_{\text{query}})&=\mathbb{E}_{l}[(f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})-\mathbb{E}[y_{\text{query}}\mid x_{\text{query}}])^{2}]\\ &=\mathbb{E}_{l}[((f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})-\mathbb{E}_{l}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})])\\ &+(\mathbb{E}_{l}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})]-\mathbb{E}[y_{\text{query}}\mid x_{\text{query}}]))^{2}]\\ &=\text{Var}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})]+\left(\text{Bias}^{2}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})]\right)+\sigma^{2}\end{split}$ (2)
LTM’s variance decreases with more in-context examples. Assume $f_{\theta}$ to be $c$-Lipschitz continuous with a constant $c=(c_{1},\ldots,c_{n})$ where each $c_{i}$ scales as $\delta i^{-\alpha}$, $\delta$ is a positive constant, and $\alpha>0.5$. Then, for two independently sampled contexts $(x_{1},y_{1},\ldots,x_{n},y_{n},x_{\text{query}})$ and $(x_{1}^{\prime},y_{1}^{\prime},\ldots,x_{n}^{\prime},y_{n}^{\prime},x_{\text{query}})$ from an unbounded $D_{stream}$, the inequality $|f_{\theta}(x_{\text{query}}\mid(x_{1},y_{1},\ldots,x_{n},y_{n}))-f_{\theta}(x_{\text{query}}\mid(x_{1}^{\prime},y_{1}^{\prime},\ldots,x_{n}^{\prime},y_{n}^{\prime}))|\leq\sum_{i=1}^{n}c_{i}\mathbbm{1}_{\{x_{i}\neq x^{\prime}_{i}\}}$ holds [97]. Applying McDiarmid’s Inequality [101], for any $t>0$, the tail probability bound is:
$Pr(|f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})-\mathbb{E}_{l}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})]|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{|c|^{2}_{2}}\right)$ (3)
where $|c|^{2}_{2}=\sum_{i=1}^{\infty}(\delta i^{-\alpha})^{2}$, converging due to $\alpha>0.5$. By the Borel-Cantelli lemma, $f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})$ converges almost surely to the expected prediction $\mathbb{E_{s}}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})]$ as $n\rightarrow\infty$ [81]. This setup implies that as more samples are added to the prompt, the model’s sensitivity to small input changes diminishes, reducing error volatility.
LTM’s biases inherited from in-context examples. Assume $f_{\theta}$ mimics gradient descent steps on examples $(x_{1},y_{1},\ldots,x_{n},y_{n},x_{\text{query}})$, the model’s average prediction over many samples $\mathbb{E}_{l}[f_{\theta}(x_{\text{query}}\mid x_{1},y_{1},\ldots,x_{n},y_{n})]$ almost surely converges to the true conditional expectation $\mathbb{E}[y_{\text{query}}\mid x_{\text{query}}]$ as $n\to\infty$, provided there is no drift in the data distribution [94]. However, if there is a shift in the distribution, $(x_{1},y_{1},\ldots,x_{n},y_{n},x_{\text{query}})$ may become misaligned with the current distribution, leading to biased predictions that reflect outdated patterns. As a result, if $(x_{1},y_{1},\ldots,x_{n},y_{n},x_{\text{query}})$ is composed of samples that no longer represent the evolving data accurately, $f_{\theta}$ ceases to be an unbiased estimator. By keeping representative samples of reoccurring concepts, while gradually incorporating data from new concepts, one can create a balanced pool of samples to created an unbiased context $(x_{1},y_{1},\ldots,x_{n},y_{n},x_{\text{query}})$.
Limited context size. Each data point in the dataset does not have the same potential, as some of the data points can be more representative or informative than others. To make the most out of in-context examples, researchers have studied a variety of data summarizing techniques, often called sketching or set cover methods, which consist of finding a small subset of data points that cover a large dataset [102]. In the context of a pretrained transformer $f_{\theta}$, the goal is to find a sketching function $s:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{n\times d}$ such that: $\mathbb{E}_{s}[f_{\theta}(x_{\text{query}}\mid s(D_{\text{stream}}))]$. This sketching function can also be viewed in terms of coreset construction. For instance, a simple yet effective summarizing policy is reservoir sampling [103], which randomly selects a sample of $k$ items, and can be viewed as choosing a diagonal sketching matrix $S\in\mathbb{R}^{n\times n}$ where the diagonal entries are 1 if the point was picked. This is arguably the most straightforward way to reduce the size of dataset, simply picking as many points as permissible uniformly at random. Though it is rare for an algorithm to produce a coreset in this straightforward fashion, it has been shown that for class-balanced and outlier-resistant problems this approach can already provide reasonable results. Indeed, one should note that the existence of coresets is trivial, as the original data set itself is in fact a coreset. The key question is the existence of small coresets where the coreset size is sublinear in the number of data points, while at the same time being polynomial in other parameters, e.g. the dimension and desired approximation error [102].
Inductive biases in context optimization. Ultimately, adaptation requires expertise and can not result only from the data. The need for a localization strategy reinforces the role of inductive biases, i.e. the set of assumptions enforced into a model to perform adaptation, putting the human expert back at the center of learning. Examples include: the smoothness assumption, where if two points reside in a high-density region are close, then so should be their corresponding outputs; the cluster assumption, where if points are in the same cluster, they are likely to be of the same concept; the manifold assumption, where high-dimensional data lies in a low-dimensional manifold; or the temporal assumption, where recent points are likely to be of the same concept. However, in any of these cases, as concepts drift and reoccur over time, abruptly or gradually, the model may rely too heavily on outdated or unrepresentative samples, leading to systematic errors in prediction. While the variance term decreases as more context examples are considered, the challenge lies in the fact that averaging over a set of $n$ samples close to a given query point, either spatially, e.g. using cosine similarity, or temporally, e.g. using recent samples, may not necessarily correspond to a similar concept.
4 Experimental study
This section evaluates the performance of a naive streaming implementation of a large tabular model (LTM) against state-of-art algorithms for data stream mining, namely: Adaptive Random Forest(ARF) [43], Streaming Random Patches (SRP) [6], Boosting-like Online Learning Ensemble (BOLE) [104], Leverage Bagging (LevBag) [105], Extremely Fast Decision Tree (EFDT) [38], and Very Fast Decision Tree (VFDT) [36]. These algorithms were implemented within the Massive Online Analysis (MOA) framework [^1] https://github.com/Waikato/moa, and tuned for each dataset using a grid search strategy, with ensembles of up to 90 components, grace periods set to 100, 400, and 1000, and tie-splitting thresholds configured to 0.01, 0.05, and 0.1 as the hyperparameter settings.
Figure 4: Special attention mark
The LTM was implemented with the publicly available first version of TabPFN  [^2] https://github.com/PriorLabs/TabPFN on a T4 GPU. This model relies on causal attention layers typical of a standard decoder-only Transformer architecture [106], making it possible to process a streaming sequence where tokens become available one at a time [95]. However, with a special attention mask, where in-context tokens only attend to each other, with no attention to the query instances, as illustrated in Figure 4. Instead, the target to the query instances are used as a label to minimize the cross-entropy loss [34]. Since tabular columns are permutation invariant, the feature orderings and scalings are shuffled for an ensembled prediction. In these experiments, 4 permutations were performed.
Figure 5: Short- and long-term memories
For stream mining, a naive localization scheme was used, combining two memory strategies, as depicted in Figure 5: a First-In-First-Out (FIFO) short-term memory that stores $M_{\text{short}}$ recent samples, and a long-term memory of size $M-M_{\text{short}}$ which stores a representative subset of recent samples from each class so that the limited in-context coreset contains richer information regarding to the whole data stream. This allows to simultaneously view the data from the current and past concepts, as the model can adapt to new knowledge while reducing the catastrophic forgetting for the past. In these experiments, $M$ was set to 1000 instances, and $M_{\text{short}}$ to 750 instances. Moreover, an initial warming period $T_{\text{warm}}$ of 100 instances was considered to collect samples into the memory, without making predictions.
Input: $D_{\text{stream}}$, $M_{\text{short}}$, $M_{\text{long}}$, $T_{\text{warm}}$, $LTM$
$S,L\leftarrow\emptyset$;
$C[y]\leftarrow 0,\forall y\in\mathcal{Y}$;
$t_{\text{warm}}\leftarrow 0$;
for each new sample $(x_{t},y_{t})\in D$ do
       if $t_{\text{warm}}>T_{\text{warm}}$ then
             $\hat{y}\leftarrow\text{learn in-context}(LTM,S\cup L)$;
             $A_{T}\leftarrow\frac{1}{T}\sum_{t=1}^{T}\mathbb{1}(\hat{y}_{t}=y_{t})$;
       end if
      $S\leftarrow S\cup\{(x_{t},y_{t})\}$;
      if $|S|>M_{\text{short}}$ then
             $(x_{\text{old}},y_{\text{old}})\leftarrow$ oldest sample in $S$;
             $S\leftarrow S\setminus\{(x_{\text{old}},y_{\text{old}})\}$;
             $L\leftarrow L\cup\{(x_{\text{old}},y_{\text{old}})\}$;
             $C[y_{\text{old}}]\leftarrow C[y_{\text{old}}]+1$;
       end if
      if $|L|>M_{\text{long}}$ then
             $y_{\max}\leftarrow\arg\max_{y\in\mathcal{Y}}C[y]$;
             $(x_{\text{old}},y_{\text{old}})\leftarrow$ oldest sample in $\{(x,y)\in L\mid y=y_{\max}\}$;
             $L\leftarrow L\setminus\{(x_{\text{old}},y_{\text{old}})\}$;
             $C[y_{\text{old}}]\leftarrow C[y_{\text{old}}]-1$;
       end if
end for
Algorithm 1 Naive context optimization
The pseudocode is given in Algorithm 1. The algorithm begins by initializing the short-term memory $S$ and long-term memory $L$, along with a class count dictionary $C[y]$ for each class $y$. The size of the long-term memory is computed as $M_{\text{long}}=M-M_{\text{short}}$. After a warming period $T_{\text{warm}}$, the algorithm enters the main loop, processing each new sample $(x_{t},y_{t})$ from the data stream $D_{\text{stream}}$. For each new sample, the algorithm first adds it to the short-term memory $S$. If the short-term memory exceeds its maximum size $M_{\text{short}}$, the oldest sample is removed and moved to the long-term memory $L$. The class count for the corresponding class of the removed sample is updated in the dictionary $C$. Once the long-term memory exceeds its capacity $M_{\text{long}}$, the algorithm identifies the most overrepresented class, $y_{\max}$, in long-term memory $L$. The oldest sample from this overrepresented class is then removed to maintain balance in the memory, ensuring that the system does not retain too many samples from any single class. Throughout the process, the algorithm continuously updates both memories while ensuring the long-term memory remains balanced across classes, and the system adapts to new data as it arrives.
@@ 4 Experimental study
In evaluating the application of any foundation model, one can consider both intrinsic and extrinsic evaluation metrics [107]. Extrinsic evaluation refers to measuring the performance of the adapted foundational model on a downstream task, whereas intrinsic evaluation refers to directly measuring the foundational model’s quality. Since this work is concerned with the adaptability of LTMs to drifting streams, only extrinsic performance evaluation is performed, using a prequential evaluation strategy, where each instance is first used to test then update the classifier in an online manner (instance by instance) [108]. The used data streams are from the USP data stream repository [^3] https://sites.google.com/view/uspdsrepository, encompassing both binary and multiclass classification tasks: NOAA Weather (18,159 instances), SmartMeter LADPU (22,950), Electricity (45,312), Rialto Bridge Timelapse (82,250), Posture (164,860), Forest CoverType (581,012) and PokerHand (829,201).
Figure 6: Nemeyi test at p-value = 5%
Table 1 and Figure 6 present a comparison of LTM’s accuracy against the aforementioned baselines. LTM consistently outperforms, with an average accuracy of 82.83% and optimal rank of 1. The top competing algorithms, ARF and SRP, achieved average accuracies of 77.47% and 78.87%, respectively. The most significant performance gains were observed on the POKER, RIALTO, METER, and NOAA datasets, where LTM outperformed the best alternative by 9.4%, 6.81%, 5.55%, and 3.94%, respectively. Moreover, it is important to highlight that LTM eliminates the need for hyperparameter tuning, as it is a pre-tuned model. In contrast, the baseline algorithms can exhibit significant variability in their results depending on the parameter settings. This issue is especially pronounced for single tree models like EFDT and HT, which show high standard deviations of up to 4.1 and 7.02, respectively. Although increasing the number of ensembles provides more stability, SRP, LevBag, BOLE, and ARF still exhibit standard deviations of up to 1.81, 5.78, 1.54, and 0.43, respectively. This underscores the advantage of LTM’s consistency in performance across varying conditions.
Table 1: Prequential accuracy
Dataset LTM ARF (n=90) BOLE (n=50) LevBag (n=45) SRP (n=80) EFDT VFDT
NOAA 81.45 78.36 ± 0.20 73.45 ± 0.51 75.51 ± 0.23 78.35 ± 0.18 72.34 ± 0.43 70.86 ± 1.56
METER 73.56 68.02 ± 0.36 67.28 ± 1.54 63.17 ± 5.78 69.69 ± 1.81 57.94 ± 2.21 52.07 ± 2.63
ELEC 91.70 87.89 ± 0.22 90.17 ± 0.98 85.86 ± 0.16 88.70 ± 0.57 77.96 ± 1.28 77.87 ± 1.29
RIALTO 80.23 67.95 ± 0.36 47.27 ± 1.54 60.99 ± 5.78 75.12 ± 1.81 55.79 ± 2.21 28.68 ± 2.63
POSTURE 59.79 59.14 ± 0.19 48.28 ± 1.09 55.44 ± 0.69 58.14 ± 0.55 48.91 ± 0.18 52.99 ± 1.05
COVER 95.51 91.75 ± 0.14 93.33 ± 0.59 85.26 ± 0.24 94.54 ± 0.20 82.38 ± 1.08 76.98 ± 2.37
POKER 97.55 89.17 ± 0.43 80.01 ± 1.54 88.52 ± 3.12 87.58 ± 0.88 77.38 ± 4.10 77.43 ± 7.02
Average 82.83 77.47 71.40 73.54 78.87 67.53 62.41
Rank 1.0 2.86 4.57 4.29 2.71 6 6.57
Figure 7 illustrates the evolution of accuracy over time for COVER, showing that LTM experiences significantly smaller performance drops at drift points compared to the baseline learners. While these baseline algorithms demonstrate rapid online convergence, LTM’s windowing approach excels in handling concept drift, maintaining a strong focus on recent instances. It is worth noting that, although LTM outperforms in drift scenarios, the competing algorithms can match LTM’s performance in long sequences of stable concepts. However, LTM remains robust, delivering consistent performance that is largely unaffected by drifts.
Figure 7: Prequential accuracy over time
@@
To further substantiate the results, Table 2 presents the reported accuracies from the original papers, including not only their own performance but also the results of other algorithms, Online Smooth Boosting (OB) [109], Online Accuracy Updated Ensemble (OAUE) [110], Dynamic Weighted Majority (DWM) [111], and Streaming Gradient Boosted Trees (SGBT) [45]. Notably, no reported algorithm outperforms LTM.
Table 2: Prequential accuracy reported in other benchmarks
Model LTM SGBT/BOLE SRP (n=10) SRP (n=100) ARF (n=100)
ELEC COVER ELEC COVER ELEC COVER ELEC COVER ELEC COVER
ARF [43] 87.89 91.75 90.62 94.72 88.71 94.69 89.67 94.97 88.54 92.32
LevBag [105] 85.86 85.26 89.71 88.13 90.16 94.86 89.51 95.1 88.53 93.08
SRP [6] 88.7 94.54 89.68 95.34 88.82 95.25 89.86 95.35
OB [109] 89.51 92.69 85.25 90.33 89.52 92.7 87.05 86.34
OAUE [110] 88.28 90.17 87.41 92.86 86.37 92.26
DWM [111] 88.52 87 87.76 88.29 87.76 88.52
SGBT [45] 88.5 94.29
BOLE [104] 90.17 93.33 90.04 90.16
LTM 91.70 95.51 91.70 95.51 91.70 95.51 91.70 95.51 91.70 95.51
Table 3: Processing time per instance
@@
While the performance of LTM is well-established, it is essential to evaluate its adherence to key data stream mining requirements [5]: (R1) processing one example at a time and inspecting it only once, (R2) using a limited amount of memory, (R3) operating within a limited time budget, and (R4) being ready to make predictions at any point. Regarding (R1), no strict constraint prevents an algorithm from temporarily storing examples internally. As long as these stored instances are eventually discarded to satisfy (R2), an algorithm remains compliant with this requirement. In terms of (R2), although LTMs may require substantial memory storage, their memory footprint remains constant, enabling them to handle data volumes far exceeding the available working memory. However, memory limitations are a physical constraint that must be balanced with adherence to (R3), i.e., ensuring feasible runtime performance. The memory in LTMs can be divided into two categories: the one allocated for running statistics of long- and short-term memory and the memory used to store the foundational model. While data stream mining algorithms typically optimize memory efficiency by maintaining running statistics that directly inform predictions, LTMs inherently separate these two types of memory. The implications of this distinction are further explored in Section 5. For (R3), LTMs must scale efficiently to any number of examples, which is ensured by its linear runtime complexity concerning the number of processed instances. Furthermore, the computational cost per example is upper-bounded by the number of features, classes and instances within the given context. A key constraint is that for a LTM to function in real-time, it must process instances as fast as they arrive, with failure to do so resulting in data loss. Although LTMs may be slower than Hoeffding tree-based models, several strategies exist to mitigate this issue, depending on the specific LTM architecture. For example, Table 3 shows the processing time per sample when using the TabPFN architecture to predict ten instances in parallel within a single forward pass. Alternatively, a hypernetwork-based LTM detaches the inference cost of the LTM from that of the prediction by the generated network, suggesting a different strategy. Further details on LTM architectures are provided in Section 5. Moreover, traditional strategies can be applied on an application-specific basis, ensuring that excessive delays do not reduce the LTM’s utility for users requiring timely predictions. For instance, instances can be offloaded to other learners, or buffered processing can be used during idle periods, especially in scenarios where absolute timing is less critical, such as when classifying large, persistent data sources. Finally, concerning (R4), while conventional stream learners continuously update their models based on evolving statistics, an LTM translates these statistics into a structured representation, incorporating a selected subset of in-context examples.
5 Future research directions

### v2

@@ 4. Context optimization
[…]
5:for each new sample $(x_{t},y_{t})\in D$ do
6:  if $t_{\text{warm}}>T_{\text{warm}}$ then
7:   $\hat{y}\leftarrow\text{learn in-context}(LTM,S\cup L)$
8:   $A_{T}\leftarrow\frac{1}{T}\sum_{t=1}^{T}\mathbb{1}(\hat{y}_{t}=y_{t})$
9:  end if
10:  $S\leftarrow S\cup\{(x_{t},y_{t})\}$
11:  if $|S|>M_{\text{short}}$ then
12:   $(x_{\text{old}},y_{\text{old}})\leftarrow$ oldest sample in $S$
13:   $S\leftarrow S\setminus\{(x_{\text{old}},y_{\text{old}})\}$
14:   $L\leftarrow L\cup\{(x_{\text{old}},y_{\text{old}})\}$
15:   $C[y_{\text{old}}]\leftarrow C[y_{\text{old}}]+1$
16:  end if
17:  if $|L|>M_{\text{long}}$ then
18:   $y_{\max}\leftarrow\arg\max_{y\in\mathcal{Y}}C[y]$
19:   $(x_{\text{old}},y_{\text{old}})\leftarrow$ oldest sample in $\{(x,y)\in L\mid y=y_{\max}\}$
20:   $L\leftarrow L\setminus\{(x_{\text{old}},y_{\text{old}})\}$
21:   $C[y_{\text{old}}]\leftarrow C[y_{\text{old}}]-1$
22:  end if
23:end for
5. Experimental study
This section evaluates the performance of the proposed context optimization of a large tabular model (LTM) against state-of-art algorithms for data stream mining, namely: Adaptive Random Forest (ARF) (gomes2017adaptive), Streaming Random Patches (SRP) (gomes2019streaming), Boosting-like Online Learning Ensemble (BOLE) (de2016boosting), Leverage Bagging (LevBag) (bifet2010leveraging), Extremely Fast Decision Tree (EFDT) (manapragada2018extremely), and Very Fast Decision Tree (VFDT) (domingos2000mining). These algorithms were implemented within the Massive Online Analysis (MOA) framework [^1] https://github.com/Waikato/moa, and tuned for each dataset using a grid search strategy, with ensembles of up to 90 components, grace periods set to 100, 400, and 1000, and tie-splitting thresholds configured to 0.01, 0.05, and 0.1 as the hyperparameter settings. The Hoeffding bound confidence level was fixed at $1\times 10^{-7}$ across all experiments. Adaptive Naive Bayes prediction was used at the leaf level for all models. For the LTM, we use the publicly available first version of the transformer-based TabPFN [^2] https://github.com/PriorLabs/TabPFN. A key practical advantage of LTMs is their minimal reliance on hyperparameter tuning. In TabPFN, the primary source of variability arises from the ensemble prediction mechanism, in which feature orderings and scalings are randomly permuted to account for the permutation invariance of tabular columns. In our experiments, 4 permutations were used. Increasing the number of permutations reduces stochasticity and yields a closer approximation to the Bayesian posterior, typically resulting in monotonic performance improvements. This behavior contrasts with tree-based ensembles, where adding models introduces explicit bias–variance–diversity trade-offs. The proposed dual-memory mechanism also introduces very few parameters. For context optimization, $M$ was set to 600, 800, and 1000 instances, and $M_{\text{short}}$ to 0.65, 0.75, 0.85 instances. Moreover, an initial warming period $T_{\text{warm}}$ of 100 instances was considered to collect samples into the memory, without making predictions.
Figure 7. Prequential accuracy over time
@@ 5. Experimental study
In evaluating the application of any foundation model, one can consider both intrinsic and extrinsic evaluation metrics (bommasani2022opportunities). Extrinsic evaluation refers to measuring the performance of the adapted foundational model on a downstream task, whereas intrinsic evaluation refers to directly measuring the foundational model’s quality. Since this work is concerned with the adaptability of LTMs to streams, only extrinsic performance evaluation is performed, using a prequential evaluation strategy, where each instance is first used to test then update the classifier in an online manner (instance by instance) (gama2013evaluating). The used data streams are from the USP data stream repository [^3] https://sites.google.com/view/uspdsrepository, encompassing both binary and multiclass classification tasks: NOAA Weather (18,159 instances), SmartMeter LADPU (22,950), Electricity (45,312), Rialto Bridge Timelapse (82,250), Posture (164,860), Forest CoverType (581,012) and PokerHand (829,201). Table 1 and Figure 8 compare LTM against all baseline algorithms. LTM achieves the highest overall performance, with an average accuracy of 82.65% and a mean rank of 1. The strongest baselines, ARF and SRP, reach 77.47% and 78.87%, respectively. The largest improvements are observed on the POKER, METER, RIALTO, and NOAA datasets, where LTM exceeds the best alternative by 8.34 pp, 3.93 pp, 3.91 pp, and 2.93 pp, respectively. Two-tailed t-tests (p << 0.05) confirm LTM’s superiority in all cases except against ARF on POSTURE, BOLE on ELEC/COVER, and SRP on COVER.
Figure 8. Nemeyi test at p-value = 5%
Grid searches over total memory sizes {600, 800, 1000} and short-term ratios {0.65, 0.75, 0.85} produce highly stable accuracies, with standard deviations up to only 0.31, with the exception of RIALTO. In contrast, several baselines exhibit substantial sensitivity to hyperparameters. Single-tree models such as EFDT and VFDT show standard deviations as high as 4.10 and 7.02, while ensemble methods such as SRP, LevBag, and BOLE can still fluctuate up to 1.81, 5.78, and 1.54, respectively. These results highlight LTM’s robustness and consistency across conditions. Table 2 reports the ablation study of each individual component of the dual-memory strategy. Incorporating the dual strategy consistently improves performance over the long-term component alone, with an average gain of +3.24 (range +2.33 to +4.59). These gains are attenuated in datasets with large number of classes and highly volatile class distributions, as reflected by the weak performance of the Majority-Class baseline. The dual strategy also improves over short-term memory alone (+0.31 on average), though gains are smaller and sometimes negative (ELEC, COVER). This is likely attributable to strong temporal dependencies, as suggested by the high accuracy of the No-Change classifier. Increasing context size from 600 to 1000 yields an average improvement of +0.75, statistically significant for most datasets.
Table 2. Ablation study (*: no statistical difference)
Dataset Dual - Long Dual - Short 1000 - 600
NOAA +4.59 +0.89 +0.50*
METER +3.35 +0.27* +0.98
ELEC +4.11 -0.22* +0.89
RIALTO +3.45 +0.91 +1.54
POSTURE +2.33 +0.24* +0.39*
COVER +2.48 -0.12* +0.46*
POKER + 2.35 +0.18* +0.47
Average +3.24 +0.31 +0.75
@@
Figure 7 illustrates the evolution of accuracy over time for COVER, showing that LTM experiences significantly smaller performance drops at drift points. While these baselines re-stabilize quickly, LTM’s windowing approach excels in handling drifts, maintaining a strong focus on recent instances. In extended intervals of stable concepts, however, the baselines tend to converge toward LTM’s accuracy. To further substantiate the results, Table 3 presents the reported accuracies from the original papers of the various baselines, including not only their own performance but also the results of other algorithms: Online Smooth Boosting (OB) (chen2012online), Online Accuracy Updated Ensemble (OAUE) (brzezinski2014combining), Dynamic Weighted Majority (DWM) (kotler2003dynamic), and Streaming Gradient Boosted Trees (SGBT) (gunasekara2024gradient). Notably, none report results outperforming LTM.
Table 3. Prequential accuracy reported in other benchmarks
Model SGBT/BOLE SRP (n=100) ARF (n=100)
ELEC COVER ELEC COVER ELEC COVER
ARF 90.62 94.72 88.71 94.69 89.67 94.97
LevBag 89.71 88.13 90.16 94.86 89.51 95.1
SRP 89.68 95.34 88.82 95.25 89.86 95.35
OB 89.51 92.69 85.25 90.33 89.52 92.7
OAUE 88.28 90.17 87.41 92.86
DWM 88.52 87 87.76 88.29 87.76 88.52
SGBT 88.5 94.29
BOLE 90.04 90.16
LTM 91.82 95.53 91.82 95.53 91.82 95.53
While the performance of LTM is well-established, it is essential to evaluate its adherence to key stream mining requirements (bifet2009adaptive; lourencco2025device): (R1) processing one example at a time and inspecting it only once, (R2) using a limited amount of memory, (R3) operating within a limited time budget, and (R4) being ready to make predictions at any point. Regarding (R1), no strict constraint prevents an algorithm from temporarily storing examples internally. As long as these stored instances are eventually discarded to satisfy (R2), an algorithm remains compliant with this requirement. In terms of (R2), although LTMs may require substantial memory storage, their memory footprint remains constant, enabling them to handle data volumes far exceeding the available working memory. In this regard, memory limitations are a physical constraint that must be balanced with adherence to (R3), i.e., ensuring feasible runtime performance. The memory in LTMs can be divided into two categories: the one allocated for running statistics of long- and short-term memory and the memory used to store the foundational model. While data stream mining algorithms typically optimize memory efficiency by maintaining running statistics that directly inform predictions, LTMs inherently separate these two types of memory. For (R3), LTMs must scale efficiently to any number of examples, which is ensured by its linear runtime complexity concerning the number of processed instances. Furthermore, the computational cost per example is upper-bounded by the number of features, classes and instances within the given context. A critical requirement for real-time operation is that the LTM processes instances as fast as they arrive, as failure to do so leads to data loss. For reference, in MOA, an ARF with 100 hoeffding trees processes 100k samples with 30 features in 2.4 ms per instance. While LTMs may be slower, several strategies can mitigate this, depending on the architecture. For example, Table 4 reports the processing time per sample when using TabPFN on a NVIDIA Tesla T4 GPU to predict ten instances in parallel in a single forward pass. Alternatively, a hypernetwork-based LTM (bonet2024hyperfast) detaches the inference cost of the LTM from that of the prediction by the generated network, suggesting a different strategy. Classic stream management techniques can also be applied on an application-specific basis to ensure timely predictions. For example, instances can be offloaded to other learners or buffered for processing during idle periods. Finally, concerning (R4), while conventional stream learners continuously update their models based on evolving statistics, a LTM translates these statistics into a structured representation, incorporating a selected subset of in-context examples.
Table 4. Processing time per instance
@@
6. Conclusions